7. Find the maximum likelihood estimate of parameter p of the binomial distribution.
Answer #1

For binomial distribution (n, p), the pmf is,

P(X = x) = inom{n}{x} p^x (1-p)^{n-x}

For random samples X1, X2, ...., Xk, the likelihood function is,

L(p) = P(X = x_1) * P(X = x_2) * ...* P(X = x_k)

n-T2 m-T1 1

t-T に! (t2

The log-likelihood function is,

log( L(p)) = y log ( n ) + yar i log(p) + y(n-xi)log(1-p)

For maximum likelihood estimate,

(n - )0

(n

Rightarrow rac{1-p}{p} = rac{sum_{i=1}^{k}(n-x_i)}{sum_{i=1}^{k}x_i }

n-Ti

Rightarrow rac{1}{p} = rac{sum_{i=1}^{k}(n-x_i)}{sum_{i=1}^{k}x_i } + 1

2 + 1

Rightarrow rac{1}{p} = rac{nk}{sum_{i=1}^{k}x_i } - 1 + 1

Rightarrow p = rac{sum_{i=1}^{k}x_i}{nk }

Rightarrow p = rac{ar{x}}{n } where ar{x}= rac{sum_{i=1}^{k}x_i}{k } (sample mean)

Thus the maximum likelihood estimate of p = ar{x} / n where ar{x} is the sample mean of the random samples.